The income elasticity of restaurant meals is 0.6. If customers income becomes 300,000 LBP instead of 200,000 LBP , the quantity demanded will
1
Expert's answer
2020-04-28T07:54:13-0400

The income elasticity of demand is positive. This means that an increase in the consumer's income will increase the quantity demanded and vice versa.


The quantity demanded will increase by 30%.


The income elasticity of demand is determined as:



"E_Y = \dfrac{\%\Delta Q}{\%\Delta Y}"

The income elasticity of demand is 0.6. Therefore:



"\dfrac{\Delta Q}{\Delta Y} = 0.6"

If the income increases from 200,000 LBP to 300,000 LBP, this is an increase in income by 50%. Therefore:



"\dfrac{\%\Delta Q}{50\%} = 0.6"

"\%\Delta Q = 0.6\times 50\%"

"\boxed{\color{red}{\%\Delta Q = 30\%}}"